OpenStudy (anonymous):

Line WX contains (-1, 2) and (4, 12) Line YZ contains points (-5, 8) and (2, -6). Lines WX and YZ are (4 points) Select one: a. perpendicular because the slopes are the same b. parallel because the product of the slopes is -1 c. perpendicular because the product of the slopes is -1 d. parallel because the slopes are the same

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

we have to find the slope of these lines using the slope formula slope = (y2 - y1) / (x2 - x1) Line WX : (-1,2) x1 = -1 and y1 = 2 (4,12) x2 = 4 and y2 = 12 now sub into the slope formula slope = (12 - 2) / (4 - (-1) slope = 10/(4 + 1) slope = 10/5 which reduces to 1/2 so the slope of line WX = 1/2 Line YZ : (5,8) x1 = 5 and y1 = 8 (-2,-6) x2 = -2 and y2 = -6 now we sub into slope formula slope = (-6-8) / (-2-5) slope = -14/-7 slope = 1/2 so the slope in line YZ = 1/2 Line WX and line YZ are : (D) parallel because the slopes are the same.

OpenStudy (anonymous):

thank you

OpenStudy (nurali):

Anytime.